OpenStudy (anonymous):

how do i graph: Q / 12-p for 010

OpenStudy (anonymous):

Choose a value of P according the information. 1st part: Q=12-P for 0<P<8 When P is 1, Q=12-1 Therefore Q=11 When P is 2, Q=12-2 Therefore Q=10 Continue putting in values of P till 7. You can make a table out of this information like so: Plotting a graph using the table gives you this:  It will be a straight line graph if drawn accurately Try using this concept for the 2nd part